OpenStudy (anonymous):

find the inverse of the given function f(x)=3/24x-13

OpenStudy (anonymous):

\[f(x)=\frac{3}{24x-13}\]?

OpenStudy (anonymous):

no its f(x)=\[\frac{ 3 }{ 24 }x-13\]

OpenStudy (anonymous):

isn't that the same as \[f(x)=\frac{1}{8}x-13\]?

OpenStudy (anonymous):

in any case, if you wanted to solve \[\frac{1}{8}x-13=5\] you would a) add 13 to get \(\frac{1}{8}x=5+13\) and then multiply by 8 to get \(x=8(5+13)\)

OpenStudy (anonymous):

that is what the inverse does, add 13, multiply by 8 i.e. \[f^{-1}(x)=8(x+13)\]

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

yw